OpenStudy (anonymous):

solve over domain -2pi

OpenStudy (campbell_st):

well start by using sin2x = 2sin(x)cos(x) so you have \[2\sin(x)\cos(x) + \sqrt{3} \cos(x) = 0\] this equation can now be factorised... to find the solutions.

OpenStudy (anonymous):

help still